OpenStudy (anonymous):

If p(1/q) = -1, then which of the following is NOT necessarily true? (a) p+q=0 (b) p-q=2p (c) pq=-p^2 (d) p^2 + q^2 = 2p^2 (e) p= -1/q

OpenStudy (goformit100):

(b) p-q=2p

OpenStudy (anonymous):

how'd u get that?

OpenStudy (zarkon):

p(1/q) = -1 so p=-q replace all the p's above with -q and see if you get equality

OpenStudy (anonymous):

i do get equality

OpenStudy (zarkon):

shouldn't on all of them..one of them is bad

OpenStudy (anonymous):

which one?

OpenStudy (zarkon):

I can't just give the answer

OpenStudy (zarkon):

notice that b) is ok p-q=2p -q-q=2(-q) -2q=-2q

OpenStudy (zarkon):

what do you get for e

OpenStudy (anonymous):

i chose that one as the answer but wasn't sure

OpenStudy (zarkon):

p= -1/q -q=-1/q this is not always true...if q=2 then -2=-1/2

OpenStudy (zarkon):

so the answer is (e)

OpenStudy (anonymous):

thank you :)